Two blocks move along a linear path on a nearly frictionless air track. One block, of mass 0.116 kg, initially moves to the right at a speed of 5.20 m/s, while the second block, of mass 0.232 kg, is initially to the left of the first block and moving to the right at 7.30 m/s. Find the final velocities of the blocks, assuming the collision is elastic.

Velocity of the .116 kg block to the right:

Velocity fo the .232 kg block to the right:
Answer #1

let
m1 = 0.232 kg

and u1 = 7.3 m/s

m2 = 0.116 kg

and u2 = 5.2 m/s

final velocitis are
Velocity fo the .232 kg block to the right is v1

Velocity of the .116 kg block to the right is v2


v1 =(m1+m2)*u1/(m1+m2) + 2*m2*u2/(m1+m2)

v1 = (0.116-0.232)*7.3/(0.116+0.232) + (2*0.116*5.2)/(0.116+0.232)


v1 = -2.43 + 3.46 = 1.03 m/s

v2 = 2*m1*u1/(m1+m2) + (m2-m1)*u2/(m1+m2)

v2 = ((2*0.232*7.3)/(0.232+0.116)) + ((0.116-0.232)*5.2/(0.116+0.232)) = 8 m/s